OpenStudy (anonymous):

Vankatlal takes money from the Employees Cooperative Society at lower rate of interest and invests in a scheme, which gives him a compound interest of 20%, compounded annually. Find the least number of complete years after which his sum will be more than double.

OpenStudy (anonymous):

x(1+r/100)^n=1/5p

OpenStudy (anonymous):

n=3 years

OpenStudy (anonymous):

@theEric

OpenStudy (shamim):

the formula is C=p(1+r)^n

OpenStudy (shamim):

given c=2p n=?

OpenStudy (anonymous):

3